OpenStudy (anonymous):

Find a polynomial equation of degree 3 such that f(0)= 31, f'(1)= 4, f"(2)= 2, and f'"(3)=6

OpenStudy (anonymous):

start with f'''(x). if f'''(3) = 6, then for simplicity, we could assign f'''(x) = 2x. the integral of f'''(x) is f''(x) + C. f''(x) = x^2 + C. let's plug in 2, and adjust C such that f''(2) = 2. f''(2) = 2^2 + C = 2. ---> C = -2 so f''(x) = x^2 - 2. is this enough info to do the rest or should i keep going?

OpenStudy (anonymous):

so is the next step x^2-2=2? i don't know how to the next step

OpenStudy (anonymous):

sorry. don't do this. it'll lead to a 4th order polynomial.

OpenStudy (anonymous):

assign f'''(x) = 6. therefore f'''(3) = 6 holds true. the integral of 6 is 6x + C = f''(x) f''(2) = 2. 6(2) + C = 2 C = -10 so f''(x) = 6x -10. now to find f'(x), we take the integral of f''(x). do you know what that is?

OpenStudy (anonymous):

x^6-10x+c?

OpenStudy (anonymous):

close. it's 3x^2 - 10x + C \[\int\limits_{}^{} x^n = \frac{ x^{n+1} }{ n+1 }\] \[\int\limits_{}^{}6x = 6\int\limits_{}^{}x = 6*\frac{ x^2 }{ 2 } = 3x^2\]

OpenStudy (anonymous):

\[\int\limits_{}^{}x^4 = \frac{ x^5 }{ 5 }\]

OpenStudy (anonymous):

i didn't even consider, have you learned integrals yet?

OpenStudy (anonymous):

there's a longer way to do it with only derivatives

OpenStudy (anonymous):

no we haven't learned integrals yet haha i was wondering why this made zero sense to me

OpenStudy (anonymous):

lol :P

OpenStudy (anonymous):

so a 3rd order polynomial looks like this:\[f(x) = ax^3 + bx^2 + cx + d\] the derivative of that, 2nd order poly, would look like this:\[f'(x) = 3ax^2 + 2bx + c\]and\[f''(x) = 6ax + 2b\]\[f'''(x) = 6a\] given our info: f'''(3) = 6 we an plug 3 in our "general" f'''(x) to find out what a is 6 = 6a a = 1. this is true for all of the equations derived from each other. given: f''(2) = 2, a = 1 and f''(x) = 6ax + 2b you can figure out b 2 = 6(1)(2) + 2b ----> b = -5 and then you can find c, then d

OpenStudy (anonymous):

i'll start the next one: given f'(1) = 4:\[4 = 3(1)(1^2) + 2(-5)(1) + c\] solve for c

OpenStudy (anonymous):

so c= 11

OpenStudy (anonymous):

and then f(x)= x^3 -5x^2 +B

OpenStudy (anonymous):

and d= 31?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay i got it :) thank you so much!!

OpenStudy (anonymous):

glad i could help :)